OpenStudy (precal):

Differential equations-need someone to double check my solution please dy/dx=ysec^2x and y=5 when x=0, then y=?

OpenStudy (precal):

I separated the terms and got (1/y) dy=sec^2(x)dx I integrate each side

OpenStudy (precal):

ln(y)=tanx+C I sub x=0 and y=5 got ln5=tan0+c ln5=c

OpenStudy (precal):

applied e as a base to both side (not sure I am describing that correctly) y=e^(tanx+ln5) y=e^tanx (e^(ln5)) y=5e^(tanx) that is my solution

OpenStudy (precal):

@amistre64

OpenStudy (ankitshaw):

you are correct...

OpenStudy (precal):

thanks

OpenStudy (amistre64):

pfft, i could have said precal was correct too .. ;)

OpenStudy (precal):

thanks